# Tutorial on amortized optimization

### v4

@@ 2.1 Defining the model $\hat{y}_{\theta}(x)$
The model $\hat{y}_{\theta}(x):{\mathcal{X}}\times\Theta\rightarrow{\mathcal{Y}}$ predicts a solution to eq. 1.1. In many applications, the best model design is an active area of research that is searching for models that are expressive and more computationally efficient than the algorithms classically used to solve the optimization problem. Section 2.1.1 starts simple with fully-amortized models that approximate the entire solution to the optimization problem with a single black-box model. Then section 2.1.2 shows how to open up the model to include more information about the optimization problem that can leverage domain knowledge with semi-amortized models.
@@ 2.1.1 Fully-amortized models
A fully-amortized model $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ maps the context to the solution of eq. 1.1 and does not access the objective $f$.
I use the prefix “fully” to emphasize that the entire computation of the solution to the optimization problem is absorbed into a black-box model that does not access the objective $f$. The prefix “fully” can be omitted when the context is clear because most amortization is fully amortized. These are standard in amortized variational inference (section 3.1) and policy leaning (section 3.6), that typically use feedforward neural networks to map from the context space ${\mathcal{X}}$ to the solution of the optimization problem living in ${\mathcal{Y}}$. Fully-amortized models are remarkable because they are often successfully able to predict the solution to the optimization problem in eq. 1.1 without ever accessing the objective of the optimization problem after being trained.
@@ 2.1.2 Semi-amortized models
A semi-amortized model $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ maps the context to the solution of the optimization problem and accesses the objective $f$ of eq. 1.1, typically iteratively.
@@
[…]
One of the most common semi-amortized model is to parameterize and integrate an optimization procedure used to solve eq. 1.1 into the model $\hat{y}_{\theta}$, such as gradient descent [Andrychowicz et al., 2016, Finn et al., 2017, Kim et al., 2018]. This optimization procedure is an internal part of the amortization model $\hat{y}_{\theta}$, often referred to as the inner-level optimization problem in the bi-level setting that arises for learning.
@@
Semi-amortized models over the domain can go significantly beyond gradient-based models and in theory, any algorithm to solve the original optimization problem in eq. 1.1 can be integrated into the model. Section 2.2.2 further discusses the learning of semi-amortized models by unrolling that are instantiated later:
  • Section 3.2 discusses how Gregor and LeCun [2010] integrate ISTA iterates [Daubechies et al., 2004, Beck and Teboulle, 2009] into a semi-amortized model.
  • Section 3.4.1 discusses models that integrate fixed-point computations into semi-amortized models. Venkataraman and Amos [2021] amortize convex cone programs by differentiating through the splitting cone solver [O’donoghue et al., 2016] and Bai et al. [2022] amortize deep equilibrium models [Bai et al., 2019, 2020].
  • Section 3.4.5 discusses RLQP by Ichnowski et al. [2021] that uses the OSQP solver [Stellato et al., 2018] inside of a semi-amortized model.
@@
Examples. The iterative gradient updates in eq. 2.1 can be replaced with a learned update function as in Ravi and Larochelle [2017], Li and Malik [2017a], Andrychowicz et al. [2016], Li and Malik [2017b]. These model the past sequence of iterates and learn how to best-predict the next iterate, pushing them towards optimality. This can be done with a recurrent cell $g$ such as an LSTM [Hochreiter and Schmidhuber, 1997] or GRU [Cho et al., 2014] and leads to updates of the form
@@
where each call to the recurrent cell $g$ takes a hidden state $z$ along with an iterate and the derivative of the objective. This endows $g$ with the capacity to learn significant updates leveraging the problem structure that a traditional optimization method would not be able to make. In theory, traditional update rules can also be fallen back on as the gradient step in eq. 2.1 is captured by removing the hidden state $z$ and setting
@@
Latent semi-amortized models are a budding topic and can excitingly learn many other latent representations that go beyond iterative gradient updates in the original space. Luo et al. [2018], Amos and Yarats [2020] learn a latent domain connected to the original domain where the latent domain captures hidden structures and redundancies present in the original high-dimensional domain ${\mathcal{Y}}$. Luo et al. [2018] consider gradient updates in the latent domain and Amos and Yarats [2020] show that the cross-entropy method [De Boer et al., 2005] can be made differentiable and learned as an alternative to gradient updates. Amos et al. [2017] unrolls and differentiates through the bundle method [Smola et al., 2007] in a convex setting as an alternative to gradient steps. The latent optimization could also be done over a learned parameter space as in POPLIN [Wang and Ba, 2020], which lifts the domain of the optimization problem eq. 1.1 from ${\mathcal{Y}}$ to the parameter space of a fully-amortized neural network. This leverages the insight that the parameter space of over-parameterized neural networks can induce easier non-convex optimization problems than in the original space, which is also studied in Hoyer et al. [2019].
@@
Starting with a single-step model. This section instantiates a single-step model similar to eq. 2.1 that parameterizes the initial iterate $\hat{y}^{0}_{\theta}(x)\vcentcolon=\theta$ and takes one gradient step:
@@
Interpreting $\hat{y}_{\theta}(x)$ as a model is non-standard in contrast to other parametric models because it makes the optimization step internally part of the model. Gradient-based optimization of losses with respect to the model’s parameters, such as eqs. 2.9 and 2.10 requires the Jacobian of $\hat{y}_{\theta}(x)$ w.r.t. the parameters, i.e. $\mathrm{D}_{\theta}[\hat{y}_{\theta}(x)]$ (or Jacobian-vector products with it). Because $\hat{y}_{\theta}(x)$ is an optimization step, the derivative of the model requires differentiating through the optimization step, which for eq. 2.4 is
@@
and requires the Hessian of the objective. In Finn et al. [2017], $\nabla_{y}^{2}f$ is the Hessian of the model’s parameters on the training loss (!) and is compute- and memory-expensive to instantiate for large models. In practice, the Hessian in eq. 2.5 is often never explicitly instantiated as optimizing the loss only requires Hessian-vector products. The Hessian-vector product can be computed exactly or estimated without fully instantiating the Hessian, similar to how computing the derivative of a neural network with backprop does not instantiate the intermediate Jacobians and only computes the Jacobian-vector product. More information about efficiently computing Hessian-vector products is available in Pearlmutter [1994], Domke [2012]. Jax’s autodiff cookbook [Bradbury et al., 2020] further describes efficient Hessian-vector products. Before discussing alternatives, the next portion derives similar results for a $K$-step model.
Multi-step models. Eq. 2.4 can be extended to the $K$-step setting with
@@
where the base $\hat{y}_{\theta}^{0}(x)\vcentcolon=\theta$ as before. Similar to eq. 2.5, the derivative of a single step is
@@
where $\mathrm{D}_{\theta}[\hat{y}_{\theta}^{0}(x)]=I$ at the base case. Computing eq. 2.8 is now $K$ times more challenging as it requires the Hessian $\nabla_{y}^{2}f$ at every iteration of the model. While using Hessian-vector products can alleviate some computational burden of this term, it often still requires significantly more operations than most other derivatives.
Computationally cheaper alternatives. The first-order MAML baseline in Finn et al. [2017] suggests to simply not use the second-order terms $\nabla_{y}^{2}f$ here, approximating the model derivative as the identity, i.e. $\mathrm{D}_{\theta}[\hat{y}_{\theta}^{K}(x)]\approx I$, and relying on only information from the outer loss to update the parameters. They use the intuition from Goodfellow et al. [2015] that neural networks are locally linear and therefore these second-order terms of $f$ are not too important. They show that this approximation works well in some cases, such as MiniImagenet [Ravi and Larochelle, 2017]. The MAML+++++ + extension by Antoniou et al. [2019] proposes to use first-order MAML during the early phases of training, but to later add back this second-order information. Nichol et al. [2018] further analyze first-order approximations to MAML and propose another approximation called Reptile that also doesn’t use this second-order information. These higher-order terms also come up when unrolling in the different bi-level optimization setting for hyper-parameter optimization, and Lorraine et al. [2020, Table 1] gives a particularly good overview of approximations to these. Furthermore, memory-efficient methods for training neural networks and recurrent models with backpropagation and unrolling such as Gruslys et al. [2016], Chen et al. [2016] can also help improve the memory utilization in amortization models.
Parameterizing and learning the objective. While this section has mostly not considered the setting when the objective $f$ is also learned, the second-order derivatives appearing in eq. 2.8 also cause issues in when the objective is parameterized and learned. In addition to learning an initial iterate, Belanger et al. [2017] learn the objective $f$ representing an energy function. They parameterize $f$ as a neural network and use softplus activation functions rather than ReLUs to ensure the objective’s second-order derivatives are non-zero.
@@ 2.1.3 Models based on differentiable optimization
As discussed in section 2.2, the model typically needs to be (sub-)differentiable with respect to the parameters to attain the Jacobian $\mathrm{D}_{\theta}[\hat{y}_{\theta}]$ (or compute Jacobian-vector products with it) necessary to optimize the loss. These derivatives are standard backprop when the model is, for example, a full-amortized neural network, but in the semi-amortized case, the model itself is often an optimization process that needs to be differentiated through. When the model updates are objective-based as in eq. 2.1 and eq. 2.2, the derivatives with respect to $\theta$ through the sequence of gradient updates in the domain can be attained by seeing the updates as a sequence of computations that are differentiated through, resulting in second-order derivatives. When more general optimization methods are used for the amortization model that may not have a closed-form solution, the tools of differentiable optimization [Domke, 2012, Gould et al., 2016, Amos and Kolter, 2017, Amos, 2019, Agrawal et al., 2019a] enable end-to-end learning.
@@ 2.2 Learning the model’s parameters $\theta$
After specifying the amortization model $\hat{y}_{\theta}$, the other major design choice is how to learn the parameters $\theta$ so that the model best-solves eq. 1.1. Learning is often a bi-level optimization problem where the outer level is the parameter learning problem for a model $\hat{y}_{\theta}(x)$ that solves the inner-level optimization problem in eq. 1.1 over the domain ${\mathcal{Y}}$. While defining the best loss is application-specific, most approaches can be roughly categorized as 1) regressing a ground-truth solution (section 2.2.1), or 2) minimizing the objective (sections 2.2.1 and 2.2.3), which fig. 2.2 illustrates. Optimizing the model parameters here can in theory be done with most parameter learning methods that incorporate zeroth-, first-, and higher-order information about the loss being optimized, and this section mostly focuses on methods where $\theta$ is learned with a first-order gradient-based method such as Nesterov [1983], Duchi et al. [2010], Zeiler [2012], Kingma and Ba [2015]. The rest of this section discusses approaches for designing the loss and optimizing the parameters with first-order methods (sections 2.2.1 and 2.2.1) when differentiation is easy or zeroth-order methods (section 2.2.3) otherwise, e.g., in non-differentiable settings.
@@ 2.2.1 Choosing the objective for learning
Regression-based learning works the best for distilling known solutions into a faster model that can be deployed at a much lower cost, but can otherwise start failing to work. In RL and control, regression-based amortization methods are referred to as behavioral cloning and is a widely-used way of recovering a policy using trajectories observed from an expert policy. Using regression is also advantageous when evaluating the objective $f(y;x)$ incurs a computationally intensive or otherwise complex procedure, such as an evaluation of the environment and dynamics in RL, or for computing the base model gradients when learning parameter optimizers. These methods work well when the ground-truth solutions are unique and semi-tractable, but can fail otherwise, i.e. if there are many possible ground-truth solutions for a context $x$ or if computing them is too intractable. After all, solving eq. 1.1 from scratch may be computationally expensive and amortization methods should improve the computation time.
@@
Eq. 2.9 can be extended to other distances defined on the domain, such as non-Euclidean distances or the likelihood of a probabilistic model that predicts a distribution of possible candidate solutions. Adler et al. [2017] propose to use the Wasserstein distance for learning to predict the solutions to inverse imaging problems.
@@
These methods use local information of the objective to provide a descent direction for the model’s parameters $\theta$. A first-order method optimizing eq. 2.10 uses updates based on the gradient
@@
where the last step is obtained by the chain rule. This has the interpretation that the model’s parameters $\theta$ are updated by combining the gradient information around the prediction $\nabla_{y}\left[f(\hat{y}_{\theta}(x);x)\right]$ shown in fig. 2.2 along with how $\theta$ impacts the model’s predictions with the derivative $\mathrm{D}_{\theta}\left[\hat{y}_{\theta}(x)\right]$. While this tutorial mostly focuses on optimizing eq. 2.11 with first-order methods that explicitly differentiate the objective, section 2.2.3 discusses alternatives to optimizing it with reinforcement learning and zeroth-order methods.
@@
A standard gradient-based optimizer for eq. 1.1 (without amortization) can be recovered from ${\mathcal{L}}_{\rm obj}$ by setting the model to the identity of the parameters, i.e. $\hat{y}_{\theta}(x)\vcentcolon=\theta$, and $p(x)$ to be a Dirac delta distribution.
This can be seen by taking $\mathrm{D}_{\theta}[\hat{y}_{\theta}(x)]=I$ in eq. 2.11, resulting in $\nabla_{\theta}{\mathcal{L}}_{\rm obj}(\hat{y}_{\theta})=\nabla_{y}f(\hat{y}_{\theta}(x);x)$. Thus optimizing $\theta$ of this parameter-identity model with gradient descent is identical to solving eq. 1.1 with gradient descent. Remark 2 shows a connection between a model trained with gradients of an objective-based loss and a non-amortized gradient-based solver for eq. 1.1. The gradient update that would originally have been applied to an iterate $y\in{\mathcal{Y}}$ of the domain is now transferred into the model’s parameters that are shared across all problem instances. This also leads to a hypothesis that objective-based amortization works best when a gradient-based optimizer is able to successfully solve eq. 1.1 from scratch. However there may be settings where a gradient-based optimizer performs poorly but an amortized optimizer excels because it is able to use information from the other problem instances.
@@
The objective-based loss in section 2.2.1 provides a starting point for amortizing with other optimality conditions or reformulations of the optimization problem. This is done when amortizing for fixed-point computations and convex optimization in section 3.4, as well as in optimal transport section 3.5.
@@
Choosing between the regression- and objective-based losses is challenging as they measure the solution quality in different ways and have different convergence and locality properties. Liu et al. [2022] experimentally compare these losses for learning to optimize with fully-amortized set-based models. Figure 2.3 illustrates that the $\ell_{2}$-regression loss (the black contours) ignores the objective values (the purple contours) and thus gives the same loss to solutions that result in significantly different objective values. This could be potentially addressed by normalizing or re-weighting the dimensions for regression to be more aware of the curvature of the objective, but this is often not done. Another idea is to combine both the objective and regression losses. Combining the losses could work especially well when only a few contexts are labeled, such as the regression and residual terms in the physics-informed neural operator paper [Li et al., 2021b]. The following summarizes some other advantages ($+$) and disadvantages (--):

### v5

@@ 2.1 Defining the model $\hat{y}_{\theta}(x)$
The model $\hat{y}_{\theta}(x):{\mathcal{X}}\times\Theta\rightarrow{\mathcal{Y}}$ predicts a solution to eq.˜1.1. In many applications, the best model design is an active area of research that is searching for models that are expressive and more computationally efficient than the algorithms classically used to solve the optimization problem. Section˜2.1.1 starts simple with fully-amortized models that approximate the entire solution to the optimization problem with a single black-box model. Then section˜2.1.2 shows how to open up the model to include more information about the optimization problem that can leverage domain knowledge with semi-amortized models.
@@ 2.1.1 Fully-amortized models
A fully-amortized model $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ maps the context to the solution of eq.˜1.1 and does not access the objective $f$.
I use the prefix “fully” to emphasize that the entire computation of the solution to the optimization problem is absorbed into a black-box model that does not access the objective $f$. The prefix “fully” can be omitted when the context is clear because most amortization is fully amortized. These are standard in amortized variational inference (section˜3.1) and policy learning (section˜3.6), that typically use feedforward neural networks to map from the context space $\mathcal{X}$} to the solution of the optimization problem living in $\mathcal{Y}$}. Fully-amortized models are remarkable because they are often successfully able to predict the solution to the optimization problem in eq.˜1.1 without ever accessing the objective of the optimization problem after being trained.
@@ 2.1.2 Semi-amortized models
A semi-amortized model $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ maps the context to the solution of the optimization problem and accesses the objective $f$ of eq.˜1.1, typically iteratively.
@@
$\hat{y}^{0}_{\theta}$$\hat{y}_{\theta}^{1}$$\ldots$$\hat{y}_{\theta}^{K}=\vcentcolon\hat{y}_{\theta}(x)$}
One of the most common semi-amortized model is to parameterize and integrate an optimization procedure used to solve eq.˜1.1 into the model $\hat{y}_{\theta}$, such as gradient descent [Andrychowicz et al., 2016, Finn et al., 2017, Kim et al., 2018]. This optimization procedure is an internal part of the amortization model $\hat{y}_{\theta}$, often referred to as the inner-level optimization problem in the bi-level setting that arises for learning.
@@
Semi-amortized models over the domain can go significantly beyond gradient-based models and in theory, any algorithm to solve the original optimization problem in eq.˜1.1 can be integrated into the model. Section˜2.2.2 further discusses the learning of semi-amortized models by unrolling that are instantiated later:
  • Section˜3.2 discusses how Gregor and LeCun [2010] integrate ISTA iterates [Daubechies et al., 2004, Beck and Teboulle, 2009] into a semi-amortized model.
  • Section˜3.4.1 discusses models that integrate fixed-point computations into semi-amortized models. Venkataraman and Amos [2021] amortize convex cone programs by differentiating through the splitting cone solver [O’donoghue et al., 2016] and Bai et al. [2022] amortize deep equilibrium models [Bai et al., 2019, 2020].
  • Section˜3.4.5 discusses RLQP by Ichnowski et al. [2021] that uses the OSQP solver [Stellato et al., 2018] inside of a semi-amortized model.
@@
Examples. The iterative gradient updates in eq.˜2.1 can be replaced with a learned update function as in Ravi and Larochelle [2017], Li and Malik [2017a], Andrychowicz et al. [2016], Li and Malik [2017b]. These model the past sequence of iterates and learn how to best-predict the next iterate, pushing them towards optimality. This can be done with a recurrent cell $g$ such as an LSTM [Hochreiter and Schmidhuber, 1997] or GRU [Cho et al., 2014] and leads to updates of the form
@@
where each call to the recurrent cell $g$ takes a hidden state $z$ along with an iterate and the derivative of the objective. This endows $g$ with the capacity to learn significant updates leveraging the problem structure that a traditional optimization method would not be able to make. In theory, traditional update rules can also be fallen back on as the gradient step in eq.˜2.1 is captured by removing the hidden state $z$ and setting
@@
Latent semi-amortized models are a budding topic and can excitingly learn many other latent representations that go beyond iterative gradient updates in the original space. Luo et al. [2018], Amos and Yarats [2020] learn a latent domain connected to the original domain where the latent domain captures hidden structures and redundancies present in the original high-dimensional domain $\mathcal{Y}$}. Luo et al. [2018] consider gradient updates in the latent domain and Amos and Yarats [2020] show that the cross-entropy method [De Boer et al., 2005] can be made differentiable and learned as an alternative to gradient updates. Amos et al. [2017] unrolls and differentiates through the bundle method [Smola et al., 2007] in a convex setting as an alternative to gradient steps. The latent optimization could also be done over a learned parameter space as in POPLIN [Wang and Ba, 2020], which lifts the domain of the optimization problem eq.˜1.1 from $\mathcal{Y}$} to the parameter space of a fully-amortized neural network. This leverages the insight that the parameter space of over-parameterized neural networks can induce easier non-convex optimization problems than in the original space, which is also studied in Hoyer et al. [2019].
@@
Starting with a single-step model. This section instantiates a single-step model similar to eq.˜2.1 that parameterizes the initial iterate $\hat{y}^{0}_{\theta}(x)\vcentcolon=\theta$ and takes one gradient step:
@@
Interpreting $\hat{y}_{\theta}(x)$ as a model is non-standard in contrast to other parametric models because it makes the optimization step internally part of the model. Gradient-based optimization of losses with respect to the model’s parameters, such as eqs.˜2.9 and 2.10 requires the Jacobian of $\hat{y}_{\theta}(x)$ w.r.t. the parameters, i.e. $\mathrm{D}_{\theta}[\hat{y}_{\theta}(x)]$ (or Jacobian-vector products with it). Because $\hat{y}_{\theta}(x)$ is an optimization step, the derivative of the model requires differentiating through the optimization step, which for eq.˜2.4 is
@@
and requires the Hessian of the objective. In Finn et al. [2017], $\nabla_{y}^{2}f$ is the Hessian of the model’s parameters on the training loss (!) and is compute- and memory-expensive to instantiate for large models. In practice, the Hessian in eq.˜2.5 is often never explicitly instantiated as optimizing the loss only requires Hessian-vector products. The Hessian-vector product can be computed exactly or estimated without fully instantiating the Hessian, similar to how computing the derivative of a neural network with backprop does not instantiate the intermediate Jacobians and only computes the Jacobian-vector product. More information about efficiently computing Hessian-vector products is available in Pearlmutter [1994], Domke [2012]. Jax’s autodiff cookbook [Bradbury et al., 2020] further describes efficient Hessian-vector products. Before discussing alternatives, the next portion derives similar results for a $K$-step model.
Multi-step models. Eq.˜2.4 can be extended to the $K$-step setting with
@@
where the base $\hat{y}_{\theta}^{0}(x)\vcentcolon=\theta$ as before. Similar to eq.˜2.5, the derivative of a single step is
@@
where $\mathrm{D}_{\theta}[\hat{y}_{\theta}^{0}(x)]=I$ at the base case. Computing eq.˜2.8 is now $K$ times more challenging as it requires the Hessian $\nabla_{y}^{2}f$ at every iteration of the model. While using Hessian-vector products can alleviate some computational burden of this term, it often still requires significantly more operations than most other derivatives.
Computationally cheaper alternatives. The first-order MAML baseline in Finn et al. [2017] suggests to simply not use the second-order terms $\nabla_{y}^{2}f$ here, approximating the model derivative as the identity, i.e. $\mathrm{D}_{\theta}[\hat{y}_{\theta}^{K}(x)]\approx I$, and relying on only information from the outer loss to update the parameters. They use the intuition from Goodfellow et al. [2015] that neural networks are locally linear and therefore these second-order terms of $f$ are not too important. They show that this approximation works well in some cases, such as MiniImagenet [Ravi and Larochelle, 2017]. The MAML++++ extension by Antoniou et al. [2019] proposes to use first-order MAML during the early phases of training, but to later add back this second-order information. Nichol et al. [2018] further analyze first-order approximations to MAML and propose another approximation called Reptile that also doesn’t use this second-order information. These higher-order terms also come up when unrolling in the different bi-level optimization setting for hyper-parameter optimization, and Lorraine et al. [2020, Table 1] gives a particularly good overview of approximations to these. Furthermore, memory-efficient methods for training neural networks and recurrent models with backpropagation and unrolling such as Gruslys et al. [2016], Chen et al. [2016] can also help improve the memory utilization in amortization models.
Parameterizing and learning the objective. While this section has mostly not considered the setting when the objective $f$ is also learned, the second-order derivatives appearing in eq.˜2.8 also cause issues in when the objective is parameterized and learned. In addition to learning an initial iterate, Belanger et al. [2017] learn the objective $f$ representing an energy function. They parameterize $f$ as a neural network and use softplus activation functions rather than ReLUs to ensure the objective’s second-order derivatives are non-zero.
@@ 2.1.3 Models based on differentiable optimization
As discussed in section˜2.2, the model typically needs to be (sub-)differentiable with respect to the parameters to attain the Jacobian $\mathrm{D}_{\theta}[\hat{y}_{\theta}]$ (or compute Jacobian-vector products with it) necessary to optimize the loss. These derivatives are standard backprop when the model is, for example, a full-amortized neural network, but in the semi-amortized case, the model itself is often an optimization process that needs to be differentiated through. When the model updates are objective-based as in eq.˜2.1 and eq.˜2.2, the derivatives with respect to $\theta$ through the sequence of gradient updates in the domain can be attained by seeing the updates as a sequence of computations that are differentiated through, resulting in second-order derivatives. When more general optimization methods are used for the amortization model that may not have a closed-form solution, the tools of differentiable optimization [Domke, 2012, Gould et al., 2016, Amos and Kolter, 2017, Amos, 2019, Agrawal et al., 2019a] enable end-to-end learning.
@@ 2.2 Learning the model’s parameters $\theta$
After specifying the amortization model $\hat{y}_{\theta}$, the other major design choice is how to learn the parameters $\theta$ so that the model best-solves eq.˜1.1. Learning is often a bi-level optimization problem where the outer level is the parameter learning problem for a model $\hat{y}_{\theta}(x)$ that solves the inner-level optimization problem in eq.˜1.1 over the domain $\mathcal{Y}$}. While defining the best loss is application-specific, most approaches can be roughly categorized as 1) regressing a ground-truth solution (section˜2.2.1), or 2) minimizing the objective (sections˜2.2.1 and 2.2.3), which fig.˜2.2 illustrates. Optimizing the model parameters here can in theory be done with most parameter learning methods that incorporate zeroth-, first-, and higher-order information about the loss being optimized, and this section mostly focuses on methods where $\theta$ is learned with a first-order gradient-based method such as Nesterov [1983], Duchi et al. [2010], Zeiler [2012], Kingma and Ba [2015]. The rest of this section discusses approaches for designing the loss and optimizing the parameters with first-order methods (sections˜2.2.1 and 2.2.1) when differentiation is easy or zeroth-order methods (section˜2.2.3) otherwise, e.g., in non-differentiable settings.
@@ 2.2.1 Choosing the objective for learning
Regression-based learning works the best for distilling known solutions into a faster model that can be deployed at a much lower cost, but can otherwise start failing to work. In RL and control, regression-based amortization methods are referred to as behavioral cloning and is a widely-used way of recovering a policy using trajectories observed from an expert policy. Using regression is also advantageous when evaluating the objective $f(y;x)$ incurs a computationally intensive or otherwise complex procedure, such as an evaluation of the environment and dynamics in RL, or for computing the base model gradients when learning parameter optimizers. These methods work well when the ground-truth solutions are unique and semi-tractable, but can fail otherwise, i.e. if there are many possible ground-truth solutions for a context $x$ or if computing them is too intractable. After all, solving eq.˜1.1 from scratch may be computationally expensive and amortization methods should improve the computation time.
@@
Eq.˜2.9 can be extended to other distances defined on the domain, such as non-Euclidean distances or the likelihood of a probabilistic model that predicts a distribution of possible candidate solutions. Adler et al. [2017] propose to use the Wasserstein distance for learning to predict the solutions to inverse imaging problems.
@@
These methods use local information of the objective to provide a descent direction for the model’s parameters $\theta$. A first-order method optimizing eq.˜2.10 uses updates based on the gradient
@@
where the last step is obtained by the chain rule. This has the interpretation that the model’s parameters $\theta$ are updated by combining the gradient information around the prediction $\nabla_{y}\left[f(\hat{y}_{\theta}(x);x)\right]$ shown in fig.˜2.2 along with how $\theta$ impacts the model’s predictions with the derivative $\mathrm{D}_{\theta}\left[\hat{y}_{\theta}(x)\right]$. While this tutorial mostly focuses on optimizing eq.˜2.11 with first-order methods that explicitly differentiate the objective, section˜2.2.3 discusses alternatives to optimizing it with reinforcement learning and zeroth-order methods.
@@
A standard gradient-based optimizer for eq.˜1.1 (without amortization) can be recovered from ${\mathcal{L}}_{\rm obj}$ by setting the model to the identity of the parameters, i.e. $\hat{y}_{\theta}(x)\vcentcolon=\theta$, and $p(x)$ to be a Dirac delta distribution.
This can be seen by taking $\mathrm{D}_{\theta}[\hat{y}_{\theta}(x)]=I$ in eq.˜2.11, resulting in $\nabla_{\theta}{\mathcal{L}}_{\rm obj}(\hat{y}_{\theta})=\nabla_{y}f(\hat{y}_{\theta}(x);x)$. Thus optimizing $\theta$ of this parameter-identity model with gradient descent is identical to solving eq.˜1.1 with gradient descent. Remark˜2 shows a connection between a model trained with gradients of an objective-based loss and a non-amortized gradient-based solver for eq.˜1.1. The gradient update that would originally have been applied to an iterate $y\in{\mathcal{Y}}$ of the domain is now transferred into the model’s parameters that are shared across all problem instances. This also leads to a hypothesis that objective-based amortization works best when a gradient-based optimizer is able to successfully solve eq.˜1.1 from scratch. However, there may be settings where a gradient-based optimizer performs poorly but an amortized optimizer excels because it is able to use information from the other problem instances.
@@
The objective-based loss in section˜2.2.1 provides a starting point for amortizing with other optimality conditions or reformulations of the optimization problem. This is done when amortizing for fixed-point computations and convex optimization in section˜3.4, as well as in optimal transport section˜3.5.
@@
Choosing between the regression- and objective-based losses is challenging as they measure the solution quality in different ways and have different convergence and locality properties. Liu et al. [2022] experimentally compare these losses for learning to optimize with fully-amortized set-based models. Figure˜2.3 illustrates that the $\ell_{2}$-regression loss (the black contours) ignores the objective values (the purple contours) and thus gives the same loss to solutions that result in significantly different objective values. This could be potentially addressed by normalizing or re-weighting the dimensions for regression to be more aware of the curvature of the objective, but this is often not done. Another idea is to combine both the objective and regression losses. Combining the losses could work especially well when only a few contexts are labeled, such as the regression and residual terms in the physics-informed neural operator paper [Li et al., 2021b]. The following summarizes some other advantages (++) and disadvantages (-):